A sample, KOH, K2CO3, dissolved in water (35cm3). Its solution was titrated with 1.058moldm-3 HCl requiring 17.7cm3 to make it colourless due to phenolphthalein solution. After adding methyl orange, the solution consumed further 6cm3 of the acid to complete the neutralisation. What is the concentration of the sample?


1
Expert's answer
2021-01-15T06:32:51-0500

"2HCl + K_2CO_3 \to 2KCl + CO_2 + H_2O\\
HCl + KOH \to KCl + H_2O"


let;

a represent hcl

b and c represent k2co3 and koh respectively


ca = 1.058 mol/dm³

va for k2co3 = 17.7cm³

va for koh = 6cm³

vb =35cm³

vc = 35cm³


Using cava/cbvb = na/nb

1.058×17.7/cb×35 = 2/1

cb = 0.268 mol/dm³


Using cava/ccvc = na/nc

1.058×6/cc×35 = 1/1

cc = 0.181 mol/dm³